### v1

# Minimizing resource overhead in fusion-based quantum computation using hybrid spin-photon devices

Abstract
We present three schemes for constructing a (2,2)-Shor-encoded 6-ring photonic resource state for fusion-based quantum computing, each relying on a different type of photon source. We benchmark these architectures by analyzing their ability to achieve the best-known loss tolerance threshold for fusion-based quantum computation using the target resource state. More precisely, we estimate their minimum hardware requirements for fault-tolerant quantum computation in terms of the number of photon sources to achieve on-demand generation of resource states with a desired generation period. Notably, we find that a group of 12 deterministic single-photon sources containing a single matter qubit degree of freedom can produce the target resource state near-deterministically by exploiting entangling gates that are repeated until success. The approach is fully modular, eliminates the need for lossy large-scale multiplexing, and reduces the overhead for resource-state generation by several orders of magnitude compared to architectures using heralded single-photon sources and probabilistic linear-optical entangling gates. Our work shows that the use of deterministic single-photon sources embedding a qubit substantially shortens the path toward fault-tolerant photonic quantum computation.
I Introduction
Fault-tolerant quantum computation (FTQC) is vital to the correct execution of the most impactful quantum algorithms, such as prime factorization [1], linear system solving [2], and chemical simulations [3]. Photonic architectures show particular promise for achieving fault-tolerance because photons are excellent carriers of quantum information. This property enables modularity, where numerous identical, smaller devices can be fabricated independently and interconnected to scale computational power. Such modularity is invaluable for improving the yield of components that meet the high quality standards necessary to satisfy the error correction thresholds [4]. Despite these advantages, the resource overheads currently required to solve industry-relevant problems remain a significant challenge [5].
In photonic quantum computation, photon loss is the dominant error. For this reason, nearly all photonic fault-tolerant architectures focus on achieving high loss thresholds [6, 7, 8, 9, 10]. One such architecture for FTQC is fusion-based quantum computation (FBQC) [6]. In this computing paradigm, which is a variation of measurement-based quantum computation [11], photonic resource states are constructed and then measured. Notably, FBQC uses finite-sized graph states that are supplied to a fusion network, where all photons are consumed by two-qubit joint measurements, or fusion gates, that serve to both construct the photonic lattice and perform the desired quantum computation on it.
FBQC is fault-tolerant provided that the supplied resource states can be produced with a per-photon loss probability well below the loss-tolerance threshold, which depends on the fusion network, the choice of resource state, and the type of fusion gates. This threshold can be increased by code concatenation, where the qubits of the resource state are themselves encoded with an inner quantum error correcting code, such as a Shor code [6, 12, 13] or a graph code [14]. This results in a bigger and more complex resource state. However, more complex states require more resources to construct that may also incur additional losses, therefore offsetting some of the original gain [15]. It is thus crucial to explore different resource-state generator (RSG) architectures to identify strategies that minimize loss while also minimizing the resource overhead.
Figure 1: Stages of fusion-based quantum computation. Single photons can be combined using linear optics and measurements to produce entangled seed states, such as a linear-cluster state or a Greenberger–Horne–Zeilinger (GHZ) state. The black lines connecting photons indicate that a control-Z operation (CZ) was applied to the two photonic qubits each initially prepared in the |+=(|$\ket{+}=(\ket{0}+\ket{1})/\sqrt{2}$ state. States which can be described using this representation are called graph states. Multiple seed states can be combined to construct a larger graph state, the target resource state, that is sent to a fusion network to implement fusion-based quantum computation. This work focuses on the stages up to and including the target state production.
In this work, we study the steps required to construct resource states for FBQC (see Figure 1). We present three different architectures for constructing the 24-photon Shor-encoded (2,2) 6-ring resource state [6] that can achieve a loss tolerance of 7.5% when used in conjunction with biased measurements [16]. The first architecture is all-photonic, requiring only a source of single photons, either heralded or deterministic. The second uses a deterministic source of caterpillar graph states [17]. The third uses a group of such sources that can be entangled using repeat-until-success (RUS) control-Z gates [18].
We begin in Section II by describing different source types and different schemes to process the photons they emit to construct the target resource state. In Section III we present the figures of merit we use to benchmark the architectures, compute their optimal bounds on resource efficiency and present simulation results. We end with a discussion in Section IV and a brief outlook in Section V.
II Architectures for Photonic Resource-State Generation
We describe four types of single-photon sources and three schemes to manipulate the emitted photons. We call the combination of a source and a scheme an RSG architecture.
Figure 2: Four different sources of quantum photonic states. a. A heralded single-photon source (SPS) constructed using intense laser pulses interacting with a non-linear material to induce four-wave mixing, whereby two pump laser photons are converted into a pair of photons of different frequencies. By filtering the light in frequency and detecting the idler photon, the presence of the single-photon signal is heralded. b. A deterministic SPS constructed from a quantum dot (QD) embedded in a micropillar cavity. The QD is excited using off-resonant laser pulses and triggered to emit a single photon at its resonance frequency, which can be separated from scattered laser photons using a highly transmittive beam splitter and filter. c. A deterministic source of caterpillar graph states constructed from a deterministic SPS containing an additional matter qubit degree of freedom such as an electron spin. The source’s qubit mediates entanglement between a chosen degree of freedom in successive photons such as polarization or time bin, producing a caterpillar graph state. d. A repeat-until-success (RUS) module where a switch is used to route photons entangled with the source’s qubit to an entangling circuit implementing a unitary transformation $U$. Afterward, the photons are measured, and, upon success, a CZ gate is applied on the two source’s qubits, thus allowing the construction of more complex graphs using fewer photons. In some case, local corrections on the source’s qubits are required, which is handled by classical feedforward communication to the lasers controlling the sources.
II.1 Photon sources
II.1.1 Heralded single-photon source
A heralded single-photon source (HSPS) is a single-photon source (SPS) that succeeds in the generation of a single photon with probability $p_{s}<1$ when triggered, and provides a heralding signal when it does succeed. This can be achieved, for example, by exploiting four-wave mixing in a material with a third-order nonlinearity [19]. Upon excitation by a strong pump laser pulse, two laser photons in the pulse may spontaneously annihilate leading to the creation of a pair of photons with a probability $p_{\textrm{pair}}$, called the idler and the signal, with different frequencies that sum to the laser frequency (see Figure 2a). The idler and signal can then be split into two different spatial modes so that the detection of the idler heralds the presence of its twin photon in the other spatial mode. Taking into account the heralding efficiency $\eta_{H}$, which is the probability that a photon is present in the signal mode when a photon was detected in the idler mode, and the idler detection efficiency $\eta_{\mathrm{det}}$, the probabilities $p_{s}$ and $p_{\textrm{pair}}$ are related by $p_{s}=\eta_{H}\eta_{\mathrm{det}}p_{\textrm{pair}}$.
Increasing the pump power improves the success probability $p_{s}$, but, when the pump is too intense, it becomes very likely that unwanted multiple pairs are produced instead. This leads to a multi-photon error probability that is proportional to $p_{s}$. Thus, $p_{s}$ is commonly limited to below $1\%$. One technique proposed to overcome this issue relies on photon-number resolving detectors to measure the idler and discard multi-photon events [20, 21]. In the limit of perfect heralding efficiency and photon-number resolution fidelity, this technique can boost the success probability up to at most $p_{s}=1/4$ [22], provided that the heralding efficiency is sufficiently high (eg. $\eta_{H}>0.99$) to reach the desired noise suppression (see Supplementary).
II.1.2 Deterministic single-photon source
A deterministic SPS (DSPS) does not suffer from a fundamental trade-off between emission probability and multi-photon errors. It produces a nearly perfect single photon every time it is triggered. It has a success probability of $p_{s}=1$, and any missing photon is attributed to the difficulty of collecting the photon. Later on, we include the photon collection efficiency in the transmission efficiency (see III.1.3).
Deterministic sources are commonly implemented using atoms [23] or artificial atoms such as quantum dots (QDs) [24]. The main challenge for this approach is designing and fabricating devices that efficiently collect the emitted photons. One very successful and robust approach is to fabricate microcavities using distributed Bragg reflectors to enhance QD emission into a single spatial mode [25, 26], that can then be collected by an optical fiber [27]. The QD emission can be triggered using excitation pulses that are detuned from the emission frequency of the QD [28, 29], and thus can be separated from the emitted single photons using a filter (see Figure 2b). Alternatively, it is possible to excite the QD resonantly, collect emission using a birefringent cavity, and filter the orthogonally-polarized photons [30].
II.1.3 Deterministic source of caterpillar graph states
Deterministic emitters provide flexibility for generating useful states. Unlike heralded sources, the system’s internal degrees of freedom can be exploited as a quantum memory to produce entangled states of light on demand through the process of sequential entanglement generation [31, 32, 33, 34]. The number of additional levels determine the complexity of the entangled state that can be produced [32, 35]. By adding just one matter qubit degree of freedom, such as an electron spin, to a two-level emitter, it is possible to generate linear cluster states and GHZ states [36, 37, 38, 39]. More generally, such an emitter can produce photonic caterpillar graph states [17], which are graph states where any dangling photons are at most one degree separated from a central one-dimensional cluster (see Figure 2c).
Caterpillar graph states are useful building blocks for constructing resource states. Since a caterpillar source can also produce GHZ states on demand, fusion gates can be boosted to a success probability of $1-1/2^{b}$, where the positive integer $b\geq 2$ is the level of boosting, while consuming $22^{b}-$ 2 additional auxiliary photons per gate [40]. Alternatively, $2b$ additional dangling photons can be added to the building block during generation to allow repeated fusion attempts without using auxiliary states, again bringing the success probability up to $1-1/2^{b}$ [41]. This latter approach requires adaptivity, which can increase the chance that remaining dangling photons in the resource state are lost. In addition, to ensure a high success probability, many dangling photons must be produced even if the first fusion attempt ends up being successful. Since the gate will succeed after two attempts on average, many unused photons must be measured to decouple them from the resource state.
II.1.4 Repeat-until-success module
To address the challenges with repeated fusions described above, an alternative is to entangle two or more caterpillar sources containing a qubit before emitting the fused graph (see Figure 2d for an example using two caterpillar sources). To entangle two sources, each source emits a photon that is sent to a small linear-optical module, where the two photons interact and get measured to implement a fusion operation. When one of the correct detection patterns is observed, this results in a controlled-Z operation applied on the spin qubits in each source. The operation is thus probabilistic, but it can be repeated until it succeeds [18]. In the absence of photon loss, the success probability $p_{s}$ can then arbitrarily approach 1 at the cost of taking more time to attempt gates.
This so-called RUS gate, and related RUS measurements, have recently been of great interest for designing hybrid spin-optical quantum computing (SPOQC) architectures tailored to quantum emitters containing a spin qubit [8, 9, 10], showing impressive thresholds. Here we demonstrate that they can also substantially improve a standard FBQC architecture, as the entanglement between the sources carries over to the emitted photons, thus allowing the generation of graph states with a more complex topology than caterpillar graphs, directly at source. Unlike repeated fusion gates, after entanglement is established, no additional dangling photons need to be produced. Therefore this approach ensures that a minimum number of photons are consumed to establish a graph edge. Moreover, the sources serve as a lossless quantum memory to store entanglement until other states are prepared.
II.2 Three RSG schemes
Figure 3: Three different schemes for constructing a Shor-encoded (2,2) 6-ring resource state. When several states need to be combined to progress to the next stage, the number is appended to the arrows. The fusion operations are boosted by generating and measuring additional auxiliary photons which are omitted for simplicity. a. An all-photonic architecture using single-photon sources that are either heralded or deterministic followed by probabilistic entanglement generation and fusion gates. b. Using a deterministic source of caterpillar graph states followed by fusion gates. c. Using a group of at least 12 deterministic caterpillar sources and repeat-until-success (RUS) CZ gates and local complementations (LC).
An RSG scheme is a sequence of steps each involving multiple quantum operations which transform an initial state into the desired target resource state. The success or failure of each step is heralded by measurement outcomes. Linking the steps together involves spatial or temporal multiplexing. A complete RSG architecture is specified by combining an RSG scheme and a source producing the initial states which the scheme takes as input.
We now propose three different such RSG schemes to construct the 24-photon Shor-encoded (2,2) 6-ring resource state originally proposed in Ref. [6]. Each scheme aims to achieve as high a performance as possible by making the most of the capabilities of each source type presented earlier.
Each scheme also makes use of a tailored multiplexing strategy to overcome probabilistic operations. This technique achieves a high generation success probability by attempting probabilistic operations many times so that a sufficient number succeed. This repetition can be done in parallel using many sources distributed in space (spatial multiplexing) [42] or using just a few sources and repeating operations sequentially in time (temporal multiplexing) [43]. Spatial multiplexing costs more physical sources, while temporal multiplexing reduces the rate at which the RSG can produce resource states. Moreover, the successfully produced states must then be routed to the desired output mode(s) in both space and time, which incurs losses.
II.2.1 All-photonic scheme
The first approach is an all-photonic scheme that aligns with the original vision of FBQC, whereby multiple unentangled single photons are passed through probabilistic entangling linear-optical circuits and fusion gates that herald the successful generation of a resource state.
Due to probabilistic operations, as the resource state becomes more complex, the success probability exponentially drops to zero. A direct approach would then cause an unmanageable exponential resource overhead. This can be alleviated by dividing up the generation procedure into successive stages where intermediate operations are heralded [15] so that complete states can be grouped together for the next operation. This dramatically increases the overall success probability and substantially decreases the amount of multiplexing needed to reach high-efficiency operation. The downside is that the number of lossy components seen by each photon, such as beam splitters and phase shifters, may increase to facilitate the selection and routing of photons between stages, possibly decreasing the maximal loss per component that the architecture can tolerate.
In Figure 3a, we present an all-photonic scheme using six multiplexing stages. We designed this particular scheme to reduce the number of states needed to attempt a new stage while also maintaining a reasonable success probability for each stage. This provides a good compromise between boosting the success probability while minimizing the number of states being routed between stages. The construction begins with single photons, produced either probabilistically or deterministically. Four photons are combined to produce an intermediate 3-photon state, called a 2-primate, which succeeds with a probability of $\sim 28.4\%$ [44]. Two of these 2-primates can then be combined into a 4-photon GHZ states with a success probability of $\sim 9.1\%$ by measuring two of the photons (see Supplementary). From there, twelve 4-GHZ states are fused together in three consecutive stages to produce the target resource state.
Normally each unboosted fusion gate is expected to succeed with a probability of $50\%$. Here, to improve the probability of successfully generating the state, we assume that each gate is boosted to reach a success probability of $75\%$ using four additional single photons (not shown in Fig. 3a) [44]. Indeed, in the low-loss regime considered in our study, the cost of producing additional photons is lower than the cost of multiplexing more fusion gates. However, boosting beyond 75% requires additional GHZ states, which are themselves costly to produce from single photons and do not justify the mild increase to success probability compared to increasing the amount of multiplexing.
This scheme can also be used starting with a source of 4-GHZ states. In that case, the first three stages of multiplexing can be neglected. We call the scheme that generates the 24-photon state directly from a source of 4-photon GHZ state a hybrid scheme.
II.2.2 Caterpillar scheme
The caterpillar scheme first applies two type-II fusion gates and three Y measurements to a 17-photon caterpillar graph to construct one-third of the resource state. The full 24-photon ring is then constructed by fusing three copies of those building blocks using three fusion gates (see Figure 3b for the fusion operations and exact caterpillar state required).
Although only a single caterpillar source is needed to produce the initial 17-photon state, the following stages require fusion gates. Thus, in this architecture, while substantially reduced, multiplexing is still required to overcome the probabilistic nature of fusion operations. As before, to optimize the success probability, all fusion gates are boosted to a success probability of $75\%$, but this time using a Bell pair (not shown in Fig. 3b) [40]. Note that a caterpillar source can also be used to efficiently boost fusion gates beyond 75%, however we find that the relative improvement compared to once-boosted gates is minor (see Supplementary) and may not justify the additional losses due to the increased complexity. Hence, we focus on the once-boosted fusion gate in our analysis.
II.2.3 RUS scheme
The last approach that we introduce involves using RUS gates to entangle a small group of caterpillar sources. There are various ways to construct the resource state using interacting emitters [45, 46, 47, 48]. Here we propose a scheme aimed primarily at minimizing the optical depth for the generated resource state. This is achieved by first creating all the necessary entanglement links between the sources and only emitting the photonic state once this first step is successful. The entanglement between the sources is directly transferred to the photons generated in this second stage. In that case, photon loss occurs only due to the source itself and a single active switch is needed to release the photon from the RUS module (Figure 2d).
Naively generating all 32 edges of the resource state requires the detection of 128 photons on average. Losing just one photon can destroy many edges and substantially increase the generation time. This problem is alleviated by recognizing that the 12 dangling photons of the resource state can be produced deterministically, provided that the highly-connected central ring of 12 qubits is first prepared across 12 sources. Furthermore, using the Graph State Compass (GSC) python package [49], we find that this smaller graph has a minimum-edge representation with just 14 edges, which can be transformed into the desired 12-qubit central ring using local complementation (LC) [^1] LC is a graph operation that corresponds to complementing the neighborhood of a vertex in a graph, and when two graphs can be obtained from one another by LC, the corresponding graph states can be obtained from one another by local Clifford operations [50]., see Figure 3c. The minimum-edge representation is hereafter called the LC-graph. This approach works only because the dangling photons can be produced afterwards, since the full 24-photon resource state is already in a minimum-edge representation.
The LC-graph has further useful properties. Most qubits have a low degree of connectivity and so photon loss is less destructive. In addition, it can be constructed by connecting together identical subgraphs in three stages. Thus, if a photon is lost when generating one subgraph, it will not destroy the others.
Once the LC-graph is complete, it is transformed into the central 12-qubit ring of the resource state using LC. Each source then emits a photon, mapping the prepared graph onto the emitted photons. Since the sources remain entangled to the photonic graph, they can emit a second photon to produce the 12 dangling photons of the final resource state. Finally each spin is measured to decouple it from the graph and reset for the next cycle.
III Architecture Efficiency Comparison
III.1 Figures of Merit
To compare architectures, we define two figures of merit quantifying the performance of an RSG: the resource efficiency $\eta_{R}$ and the transmission efficiency $\eta_{T}$. These metrics are designed to describe RSG architectures that operate on a periodic clock cycle with rate $r$, where a single resource state is produced by the RSG at the end of every cycle with a success probability $p_{\text{rsg}}$. This requirement is important to ensure synchronization between different RSGs and synchronization with the clock cycle of the fusion network required to perform FBQC.
III.1.1 Resource efficiency
To benchmark RSG overhead, a platform-agnostic metric is needed that captures trade-offs across RSG architectures. This metric could encompass factors like the energetic cost of operation, the fabrication cost, or the spatial footprint of the device. Designing an inclusive and comprehensive definition of resource efficiency is a challenge that should be addressed in a future work. Here, we consider a simple definition that enables the comparison of RSG architectures exploiting spatial and temporal multiplexing schemes.
To quantify the resource efficiency in the spatial domain, we use the ratio of the number of photons in the final resource state $N$ to the number of sources $N_{0}$ used to produce at least one such resource state with a probability greater than a predefined value. In the temporal domain, we consider the ratio of the RSG clock rate $r$ to the rate $r_{0}$ of the laser pump used trigger the sources (the internal clock of the RSG). Combining these, we define a spatio-temporal resource efficiency of:
$\eta_{R}=\frac{Nr}{N_{0}r_{0}}.$ (1)
By conservation of energy, we must always have $\eta_{R}\leq 1$. For example, one source ($N_{0}=1$) producing a large resource state of $N$ photons implies that $r\leq r_{0}/N$ as the source requires a minimum time to emit all $N$ photons. The ratio $\tau=r_{0}/r$ also quantifies the number of internal RSG clock cycles that the $N_{0}$ sources are operated for in order to attempt to produce a single resource state.
Crucially, the resource efficiency as defined above does not discriminate between spatial and temporal multiplexing. As such, it also does not capture important considerations regarding the range of values for $r$ and $r_{0}$ that are reasonable for a given material platform or fault-tolerant architecture. Thus, it is also important to interpret $\eta_{R}$ in the context of a specific implementation. We will touch on some of these considerations in Section IV.
III.1.2 Upper bound on $\eta_{R}$
The maximum possible resource efficiency, $\tilde{\eta}_{R}$, is fixed by the RSG architecture. This bound is reached in the limit of perfectly resource-efficient multiplexing [44, 51]. In this limit, an optimal multiplexer (MUX) is one that does not waste any photons. Every photon that is produced by a source will be consumed by the RSG to construct a resource state or released as part of the resource state [^2] A non-optimal MUX may instead for example only be able to route one state if two are produced at the same time.. Any excess photonic states produced during one RSG clock cycle may also be delayed to later cycles to help fill any deficit.
A spatial MUX does not decrease RSG rate ($r=r_{0}$) and so $\eta_{R}=N/N_{0}$ depends only on the number of sources $N_{0}$. If the architecture must consume an average of $N_{\text{avg}}$ photons to produce one resource state, then a necessary condition to arbitrarily improve the RSG success probability $p_{\text{rsg}}$ toward $1$ is that the $N_{0}$ sources produce a minimum of $N_{\text{avg}}$ photons on average. Thus we have $p_{s}N_{0}\geq N_{\text{avg}}$ which implies $\eta_{R}\leq p_{s}N/N_{\text{avg}}=\tilde{\eta}_{R}$. This upper bound can be generalized to $r\neq r_{0}$ by noting that an optimal temporal MUX allows one to reduce $N_{0}$ at the cost of reducing $r$ by the same proportion. Hence:
$\eta_{R}=\frac{Nr}{N_{0}r_{0}}\leq\frac{p_{s}N}{N_{\text{avg}}}=\tilde{\eta}_{R}.$ (2)
Knowing the upper bound, it is convenient to define the discrepancy between $\eta_{R}$ and $\tilde{\eta}_{R}$, which is directly attributed to the ability for a MUX to use the available photons. This leads to a natural definition of MUX efficiency $\eta_{\text{mux}}=N_{\text{avg}}r/N_{0}r_{0}$, so that
$\eta_{R}=p_{s}\eta_{\text{mux}}\frac{N}{N_{\text{avg}}}.$ (3)
This final expression is convenient as it clearly separates three independent contributions to the resource efficiency: the success probability $p_{s}$ that depends only on the source, the MUX efficiency $\eta_{\text{mux}}$ that depends only on the multiplexing network, and the efficiency of photon consumption $N/N_{\text{avg}}$ that depends only on the scheme.
The above derivation of $\eta_{\text{mux}}$ assumes that sources operate continuously for the entire RSG cycle. This is not the case for the RUS module because sources may idle after generating an edge. For example, in Figure 3c, the last operation before LC will be a RUS gate to establish the edge connecting just two of the 12 sources. When this RUS gate is being applied, the 10 other sources must wait without emitting photons, which increases the number of internal cycles per source. Thus $N_{\text{avg}}$ will underestimate the cost in time and overestimate the upper bound on resource efficiency. In this case, the upper bound can instead be computed knowing the average number of internal clock cycles $\tau_{\text{avg}}$ used to produce a resource state:
$\eta_{R}\leq\tilde{\eta}_{R}=\frac{N}{N_{0}\tau_{\text{avg}}}.$ (4)
III.1.3 Transmission efficiency
In practice, reaching the upper bound on resource efficiency is challenging, as improving $\eta_{\text{mux}}$ to unity often demands a complex multiplexing network capable of flexibly redistributing photons in space and time for use by the RSG at the right place and time [51]. However, increasing the complexity of the multiplexing network reduces the transmission efficiency $\eta_{T}$ of photons in the RSG due to the greater number of optical components, such as beam splitters, phase shifters, mode crossings, and delay lines, required for photon routing. For this reason, it is useful to distinguish between the transmission efficiency that depends on the specific choice of multiplexing network, denoted here as $\eta_{x}$, and the efficiency of photons produced by the source, which we denote $\beta$. This latter efficiency impacts all types of sources and it quantifies the transmission between the source and the input of the RSG network, including the collection efficiency of the photons at the output of the source and loss due to filtering or frequency separation. This gives a total RSG transmission efficiency of
$\eta_{T}=\beta\eta_{x}.$ (5)
Note that both $\eta_{T}$ and the RSG success probability $p_{\text{rsg}}$ impact the logical error rate of the fusion network but they do so in different ways. A non-unity $\eta_{T}$ will largely cause uncorrelated photon loss errors whereas non-unity $p_{\text{rsg}}$ results from the failure of at least one probabilistic operation inside the RSG that may introduce correlated errors. The error-correction threshold for $\eta_{T}$ depends on the resource state being produced, and is known to be $92.5\%$ [16]. However, the error-correction threshold for $p_{\text{rsg}}$ depends on how the state is constructed and would require a detailed error correction simulation of the fusion network to compute. Such a task goes beyond the scope of this comparative study and so we instead fix $p=0.999\gg 0.925$ so that any errors are dominated by independent photon loss. In addition, we verify that changing the value of $p_{\text{rsg}}$ does not significantly impact the conclusions drawn from our study.
III.1.4 Per-component maximal loss
To achieve fault-tolerance, it is necessary to have $\eta_{T}\geq 0.925$. It is then instructive to find the minimum values of $\beta$ and the transmission of basic optical components contributing to $\eta_{x}$ that are needed to reach this target.
Note that, if $\eta_{x}=1$, then we must have $\beta\geq 0.925$. Since the photon collection efficiency is characterized regularly in the literature [26, 25, 28, 52, 30, 53, 54], we do not break down the various components contributing to it. Instead, we define it as the probability of having a photon in a waveguide or fiber at the point just after any necessary filters (recall Figure 2).
Setting aside $\beta$ for now, we analyse $\eta_{x}\geq 0.925$. This requirement is satisfied by many possible configurations of the RSG that each correspond to a different value of $\eta_{R}$. It is thus important and nontrivial to find the maximum resource efficiency such that the RSG operates in the loss-tolerant regime. For a given architecture, and considering only $D$ optical components that dominate the losses with a similar loss-per component of $\sim x$, we can write $\eta_{x}=(1-x)^{D}$. Here, we denote $D$ as the optical depth of the network, or the number of the most-lossy components seen by a photon passing through the RSG. Importantly a lower bound on $D$ can be estimated for a particular architecture, which gives a per-component maximal loss $x$ such that $\eta_{x}\geq 0.925$ using:
$x\leq 1-0.925^{1/D}.$ (6)
III.2 Performance bounds
Source
Generation
scheme
MUX stages
$N_{\text{mux}}$
Optical depth
$D$
Resource
efficiency $\tilde{\eta}_{R}$
Number of SPS
$N_{0}$
Maximal loss per
component $x$
HSPS, $p_{s}$=0.005 All-photonic $6$ $36$ $0.00057\%$ $423400$ $0.22\%$
HSPS, $p_{s}$=0.05 All-photonic $6$ $36$ $0.0057\%$ $42340$ $0.22\%$
HSPS, $p_{s}$=0.25 All-photonic $6$ $36$ $0.029\%$ $8468$ $0.22\%$
DSPS All-photonic $5$ $30$ $0.11\%$ $2117$ $0.26\%$
4-GHZ source Hybrid $3$ $18$ $0.77\%$ $310$ $0.43\%$
Caterpillar source Caterpillar $2$ $12$ $8.6\%$ $28$ $0.65\%$
RUS Module RUS 00 $1$ $20\%$ $12$ $7.5\%$
Table 1: Performance bounds to generate a Shor-encoded (2,2) 6-ring resource state with the proposed architectures. For each source and generation scheme, the third column gives the number of multiplexing stages and the fourth column gives the estimated minimum optical depth associated to the multiplexing network. The maximum resource efficiency in the fifth column is given by Eq. (2), and an example of number of SPS reaching this bound is computed via Eq. 1 for a rate $r=100$MHz and displayed in the sixth column. The last column gives the maximal loss per component that each architecture can tolerate to reach a total transmission efficiency of $92.5\%$ [16] when $\beta=1$, given by Eq. 6. Since $\beta$ contributes only once to the total transmission, this maximal loss per component also estimates the required source efficiency to be $\beta\geq 1-x$. The non-RUS architectures use fusion gates boosted to a success probability of $3/4$. All sources are taken to operate at a base repetition rate of $r_{0}=1$ GHz, which is expected to be limited by recovery times for detectors and feedforward operations.
Using the above-defined metrics along with the assumption of perfectly efficient multiplexing ($\eta_{\text{mux}}=1$), we benchmark the bounds on performance for different sources combined with the schemes described in Section II.2:
  1. 1.
    1. (a)
      all-photonic scheme with a HSPS at $p_{s}=0.5\%$ representing current technology,
    2. (b)
      all-photonic scheme with a HSPS at $p_{s}=5\%$ representing realistic improvements,
    3. (c)
      all-photonic scheme with a HSPS at $p_{s}=25\%$ representing the most optimistic scenario,
    4. (d)
      all-photonic scheme with a DSPS,
    5. (e)
      a hybrid scheme that uses a deterministic source of 4-GHZ states to bypass the first steps in the all-photonic scheme,
  2. 2.
    the caterpillar source and scheme described in Section II.2.2, and
  3. 3.
    the RUS module and scheme described in Section II.2.3.
The results are summarized in Table 1.
III.2.1 All-photonic scheme
The upper bound on resource efficiency $\tilde{\eta}_{R}$ for the all-photonic scheme can be computed from Eq. 2 by first computing the average number of photons $N_{\text{avg}}$ consumed to produce one resource state on average. This can be computed following the method in Ref. [44] by knowing the success probability $p_{i}$ of each stage $i=1,2,...$ of the six-stage scheme, along with the number of auxiliary single photons $a_{i}$ used to boost fusion gates and the number of copies $c_{i}$ of the state required for the next stage.
For example, to produce on average one 2-primate we need $c_{1}=4$ single photons per attempt and we must take on average $1/p_{2}\simeq 3.5$ attempts where $p_{2}=20\sqrt{2}-28\simeq 28.4\%$ [44]. Hence, a 2-primate state requires on average $3.518\times 4\simeq 14.072$ photons to produce. The 4-GHZ state then requires $c_{2}=2$ 2-primate states costing a total of 28 photons on average, and it succeeds with a probability of $p_{3}=(3+2\sqrt{2})/64\simeq 9.11\%$ corresponding to an average of $1/p_{3}\simeq 10.98$ attempts. Thus, the average number of photons needed to produce one 4-GHZ state on average is $\sim 309$, as originally derived in Ref. [44].
In the Supplementary, we formally extend this method for any number of MUX stages. This gives a recursive formula to compute the cumulative average number of photons consumed to succeed the $i$th stage:
$N_{\text{avg},i}=c_{i}\frac{N_{\text{avg},i-1}+a_{i}}{p_{i}},$ (7)
where $N_{\text{avg},0}=1$. Note that this expression gives $N_{\text{avg},3}\simeq 2\times 309$ for producing two 4-GHZ states.
The values of $p_{i}$, $c_{i}$ and $a_{i}$ for the six-stage all-photonic scheme are summarized in Table 2. Following Eq. (7) using the values in the table, we arrive at the total of $N_{\text{avg}}=21170$ single photons for one resource state. Finally, from Eq. (1), we have the upper bound
$\tilde{\eta}_{R}=\frac{24p_{s}}{21170}$ (8)
For a heralded SPS with $p_{s}=0.5\%$ success probability, the resource efficiency is then bounded by $\eta_{R}\leq 0.00057\%$, while for a deterministic source it is $\eta_{R}\leq 0.11\%$. Assuming that the internal clock cycle (the pump repetition rate) is $r_{0}=1$ GHz and we wish to produce resource states at the rate $r=100$ MHz, these resource efficiencies correspond to a minimum number of sources of $423400$ and $2117$, respectively.
Stage, $i$ 1 2 3 4 5 6
Probability, $p_{i}$ $1$ $2820\sqrt{2}-$ 28 $\dfrac{3+2\sqrt{2}}{64}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$
Copies, $c_{i}$ 4 2 2 3 2 1
Auxiliaries, $a_{i}$ 0 0 0 8 8 8
Detected, $d_{i}$ 1 1 2 10 12 12
Cumul. $N_{\text{avg},i}$ $4$ $28$ $618$ $3339$ $11900$ $21170$
Cumulative $D_{i}$ $9$ $15$ $21$ $27$ $33$ $36$
Table 2: A summary of the relevant quantities describing the multiplexing stages of the proposed all-photonic scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary single photons needed for boosting fusion gates, $d_{i}$ is the number of detected photons, and  $N_{\text{avg, $i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
To estimate the maximal loss per component $x$ required to remain in the fault-tolerant regime, we first make an optimistic approximation that the optical depth of a photonic MUX does not depend on the number of photonic states being routed between each stage. That is, we assume that the optical depth of the RSG depends only on the type of MUX networks used and the number of MUX stages. For the type of spatial MUX network, we consider a Spanke network that uses two layers of generalized Mach-Zehnder interferometers (GMZIs) with a mode routing network in between [51]. Each GMZI is composed of many beam splitters, but each photon only sees one phase shifter using the most optimistic design [51]. Since a phase shifter is expected to be the most lossy component for a GMZI, we estimate the minimum optical depth to be $D_{\text{GMZI}}=1$. The mode routing network requires photons to be redistributed in space, which requires long propagation distances and for many lossy mode crossings. For this reason, the practical solution is to couple photons into low-loss fibers to redistribute them arbitrarily with minimal propagation loss. However, each chip-to-fiber interface causes non-negligible loss, which we identify as the most lossy component. Since it is necessary to couple out of the first GMZI and into the second GMZI, we estimate the minimum constant optical depth of the mode routing network to be $22D_{\text{coupl.}}=$ 2.
To route photons through the network, it is also necessary to reconfigure it in between each internal clock cycle. During this time, photons must be stored in a low-loss memory. Assuming that the cycle time is at least 1 ns to allow for classical processing, and knowing that state-of-the-art propagation loss in silicon nitride (SiN) is around 0.4 dB/m, the loss from a single on-chip memory is at least 60 mdB (or around 1.4% loss). Given that chip-to-fiber couplings can already be less than 60 mdB [54], a fiber memory is more practical and allows for delays much longer than 1 ns without significant loss. Thus, we also assume that every MUX stage begins and ends with a chip-to-fiber interface, bringing the minimum optical depth per spatial MUX stage to $D_{\text{s}}=2D_{\text{GMZI}}+4D_{\text{coupl.}}=6$.
To allow for temporal multiplexing, we consider the de Bruijn-based spatio-temporal network introduced in [51], which is similar to the Spanke network but includes an additional GMZI and a layer of delay lines to synchronize photons in time. In this case, the delay lines must store photons longer than 1 ns so that they can be redistributed over many internal clock cycles, leaving no choice but to implement them using fiber memories. Hence, we consider a spatio-temporal MUX to include an additional optical depth of $D_{\text{GMZI}}+2D_{\text{coupl.}}=3$ for a total of $D_{\text{b}}=9$.
Spatio-temporal multiplexing is a useful technique to reduce the number of sources while also allowing control over the RSG rate $r$ independent of the internal rate $r_{0}$. To gain this flexibility, it is only necessary to use it for one stage while the remaining stages only use a spatial MUX strategy. Furthermore, the final stage of multiplexing requires only a $1\times n$ MUX, which can be accomplished in depth 3. This brings us to our final lower-bound estimate on the optical depth of an RSG:
$D=D_{b}+(N_{\text{mux}}-2)D_{s}+3$ (9)
where $N_{\text{mux}}>1$ is the number of multiplexing stages for the scheme. Thus for the full six-stage scheme, the minimum optical depth is $D=36$, leading to a per-component maximal loss of $x=1-0.925^{1/36}=0.22\%$.
III.2.2 Caterpillar scheme
Stage, $i$ 1 2 3
Probability, $p_{i}$ 1 $\dfrac{9}{16}$ $\dfrac{27}{64}$
Copies, $c_{i}$ 1 3 1
Auxiliaries, $a_{i}$ 0 4 6
Detected, $d_{i}$ 0 11 12
Cumulative $N_{\text{avg},i}$ $17$ $112$ $280$
 Cumulative $D_{i}$ 0 9 12
Table 3: A summary of the relevant quantities describing the multiplexing stages of the proposed caterpillar scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary single photons needed for boosting fusion gates, $d_{i}$ is the number of detected photons, and $N_{\text{avg,$i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
For caterpillar scheme, we can compute the average number of photons $N_{\text{avg}}$ required to construct a single resource state using the same formula as for the all-photonic scheme. The probabilities, number of copies, and auxiliary photons used for fusion gates are all summarized in Table 3. Applying Eq. 7, we find that the average number of photons consumed to produce one resource state on average is $N_{\text{avg}}=280$. This corresponds to an upper bound on the resource efficiency of $\tilde{\eta}_{R}=8.6\%$ for the $N=24$-photon state.
For the minimum optical depth, we assume the architecture uses the same MUX networks as for the all-photonic architectures, but now using $N_{\text{mux}}=2$. This gives a minimum optical depth of $D=12$ for the caterpillar scheme, including one spatial-temporal MUX with depth 9 and one final spatial MUX with depth 3. This leads to a per-component maximal loss of $x=0.65\%$.
III.2.3 RUS scheme
To estimate the performance bounds for the RUS module architecture, we first assume that exactly $N_{0}=12$ sources are used so that the probabilistic nature of the CZ gates is only overcome using RUS, which is a form of temporal multiplexing, rather than photonic spatial multiplexing. Then, we count the average number of gate attempts made per source to estimate $\tau_{\text{avg}}$ for the full scheme, which then gives the upper bound using Eq. (4).
Each gate attempt succeeds with probability $1/2$ [18, 8, 9], therefore each round of RUS gates requires 2 attempts on average. Since the maximum degree of the LC-graph is 4 (recall Figure 3c), 4 rounds of RUS gates are necessary to produce the spin graph. Thus, all 12 sources must wait for an average of 8 RUS attempts before the LC-graph is complete. Given that local gates do not require feedforward, we assume they are completed much faster than the RUS gates so that they have a negligible impact on $r$. Finally, we account for two emission cycles that we assume operate at the same rate $r_{0}$ for simplicity. In total, the resource state can be generated using on average $\tau_{\text{avg}}=10$ internal clock cycles. This gives the RUS module approach a maximum resource efficiency of $\tilde{\eta}_{R}=24/(12\times 10)=20\%$.
As discussed in Section II.2.3, the RUS scheme was designed to minimize the optical depth to a single $1\times 2$ switch. In addition, although some platforms may use chip-to-fiber coupling to route photons through the switch, this is not strictly necessary because photons are stored in the spin memory until the switch is reconfigured. A single switch is also significantly more shallow than a GMZI. For these reasons, we estimate a minimum optical depth of $D=1$ for the RUS scheme. This implies that the per-component maximal loss is $x=7.5\%$, which is significantly higher than all other architectures thanks to the lossless memory of the RUS module. Note that, since $7.5\%$ is also substantially higher than the loss from a $1\times 2$ switch, the loss will be dominated by the source itself, meaning that a source with $\beta\geq 92.5\%$ reaches the loss-tolerant regime for this architecture.
III.3 Implementations
a.                                                            b.
Refer to caption Refer to caption
Figure 4: Resource efficiency and maximal loss that each architecture can tolerate. a. The regimes where fault-tolerant quantum computing is possible using the 24-photon Shor-encoded (2, 2) 6-ring resource state are indicated by shaded areas. The square points represent the hypothetical optimal implementations presented in Table 1. The curves illustrate the simulated trade-off between resource efficiency and maximal loss per component when using resource-sharing MUX strategies. For the HSPS, we assume a photon generation success probability of $p_{s}=5\%$, representing realistic improvements over current technology. For the RUS scheme, the solid line represents a single 12-source RSG while the dashed line represents the maximum resource efficiency achieved when using spatial multiplexing to reach $r=r_{0}$. b. The number of sources needed to reach an RSG clock rate of $r=1$ GHZ assuming an internal clock rate of $r_{0}=1$ GHZ and a maximal loss per component of $x=0.1\%$ for the best (RUS) and worst (HSPS) architectures shown in panel a. In both plots, we define the hypothetical ideal RSG architecture represented by the star point as an RSG that is able to produce resource states with perfect resource efficiency ($\eta_{R}=1$) and without photons passing through any optical components after being collected from source ($\eta_{x}=1$) so that the loss is purely determined by the source collection efficiency $\beta$.
The performance bounds computed in the previous section already provide an important order-of-magnitude comparison between the different approaches. However, as discussed in Section III.1.2, it is not straightforward to design an RSG architecture that approaches the maximum resource efficiency without also increasing optical depth, but there are some MUX strategies that help mitigate this trade-off.
Duplicate multiplexing can improve the trade-off by producing more than one copy of an intermediate state from the same register of sources to feed a single operation in the next stage. For instance, producing a 2-primate requires four identical photons. A single register of heralded sources can ensure that at least four photons are generated per clock cycle. This additional routing can be accomplished using the Spanke network discussed previously. However, in reality, the optical depth of the network will grow with an increasing amount of duplicate outputs.
Resource sharing [51] is an extension of duplicate multiplexing where copies are grouped to feed multiple operations in the next stage. For example, a register of heralded sources can be connected to two modules, each producing a 2-primate state. If four to eight photons are produced, one 2-primate state can be attempted, while eight or more allows for two attempts. This strategy reduces waste from multiplexing and improves resource efficiency toward the architecture’s upper bound. Again, the multiplexing networks described earlier give appropriate flexibility for this task. But, as with duplicate multiplexing, the cost is having a more complicated switching network that will increase the optical depth with the number of photons being routed and thus decrease the maximal loss per component that is compatible with fault-tolerance.
To illustrate these trade-offs and provide a more practical comparison of the schemes, we simulate the resource efficiency while incorporating resource-sharing strategies.
To simulate photonic multiplexing used by the all-photonic and caterpillar schemes, we assume that the RSG is composed of a series of independent RSG modules that each use maximal resource sharing at all but the last stage. These independent RSGs are then multiplexed at the final stage to achieve the desired success probability of $p_{s}=99.9\%$. For simplicity, we assume this final multiplexing stage can be accomplished in constant depth $3$ as before. The number of sources within each independent module and the resulting success probability is determined by the remaining available optical depth $D$ so that Eq. (6) is satisfied. For each MUX inside each module, we assume an optimistic sorting network of depth $\lceil\log_{2}(n)\rceil$ where $n$ is the number of states routed at the given stage. More details about the different figures and the optimization procedure are given in Supplementary. Note that less optimal sorting might be achieved with a smaller depth, but this would not allow for maximal resource sharing.
For the RUS scheme, we performed a Monte Carlo simulation of the construction procedure that is illustrated in Figure 3c (see Supplementary for details). For this scheme, we also consider two types of resource sharing. The first type is a temporal resource sharing technique that naturally occurs due to the spin memory. If a graph was not complete in time for the RSG cycle, the partially completed graph will be carried over into the next cycle, greatly increasing the probability of success. The second type is a spatial resource sharing technique that is possible due to having identical subgraphs in each stage of construction. In this case, the next stage of graph construction can commence as soon as any two subgraphs are complete, provided that the RUS gate connectivity can be reconfigured. To allow for this additional connectivity, and knowing that no delay lines are required to reconfigure the network due to the quantum memory, we assume a constant optical depth 3 on the RUS gate. Note that this does not increase the optical depth impacting the final photonic resource state, which remains $D=1$ due to the single switch.
Finally, we also consider using spatial multiplexing for the RUS scheme, where more than 12 sources are used to increase the success probability and RSG rate for the RUS scheme. The cost of this is to add an additional GMZI at the output of the RUS module to route photons into a predetermined spatial mode, although no delay lines are required.
In Figure 4a, we show how four architectures compare in this practical setting using resource sharing techniques. Notably, in this practical setting, all four architectures fail to simultaneously reach the bounds on resource efficiency and maximal loss per component presented in Table 1. However, the discrepancy is much greater for the all-photonic architectures compared to the caterpillar and RUS schemes. There is also a substantial drop in resource efficiency for the all-photonic architectures beyond a certain value of loss, which is caused by a rapid reduction in resource sharing in order to reduce the optical depth of each MUX stage to remain below the loss threshold of the resource state. On the other hand, the caterpillar scheme does not suffer a large trade off because only one MUX stage uses resource sharing and it is already very shallow due to the relatively high success probability of performing two boosted fusion gates.
It is notable that the RUS scheme also suffers from a trade-off between the resource efficiency and the maximal loss per component, with a marked decrease in resource efficiency as the maximal loss per component approaches the upper bound. This is due to a reduced efficiency of the RUS gate, causing the module to restart some operations whenever a photon is lost. This increases the amount of internal clock cycles needed to ensure $p_{\text{rsg}}$ reaches 99.9%. However, when using spatial multiplexing, the resource efficiency can approach very close to the 20% upper bound for the scheme even in the presence of significant loss. This latter result is a consequence of the fact that, for this scheme uniquely, resource sharing and multiplexing primarily impact photons used to entangle sources, rather than the final resource state itself.
Our results also show that, for a maximal loss per component of $x=99.9\%$, a multiplexed RUS module requires 6 orders of magnitude fewer sources to achieve the same rate of generation as an HSPS with $p_{s}=5\%$ (see Figure 4b). Improving $p_{s}$ to at most $25\%$ can reduce the resource overhead of the all-photonic architecture by factor of around 5, but this surprisingly provides very little improvement to the loss tolerance. Moreover, due to the shallow optical depth, the RUS scheme can achieve this resource advantage using sources with an order of magnitude higher collection loss.
IV Discussion
Natural constraints on RSG rates.
Interpreting $\eta_{R}$ in the context of a physical platform is important to ensure physical suitability. When using an RSG architecture to perform fault-tolerant computations, the rate $r_{0}$ is constrained by classical electronics due to feedforward operations and detector reset times (estimated at most $\sim$1 GHz for dedicated electronics) while the RSG rate $r$ is limited by the error correction cycle, which is dominated by the decoding time (estimated at $\sim$1 MHz [55]). If an RSG outpaces decoding, interleaving can be used to reduce the number of RSGs needed for the overall FTQC architecture [56, 5]. Thus, like temporal multiplexing, this is a strategy that can be used to trade rate to reduce the physical source overhead. However, interleaving cannot be used to increase the effective RSG efficiency above the upper bound, and it also reduces the acceptable maximal loss per component due to memories and routing.
Implication for industrial-scale applications.
Using low-loss fibers and interleaving, it is estimated that one RSG can effectively serve as several logical qubits in a surface code, leading to industrial-scale applications requiring between $10^{3}$ and $10^{5}$ RSGs [5]. Thus, with just 12 sources per RSG, an FBQC architecture using RUS-based RSGs could reach industrial-scale fault-tolerant computations using as few as 12000 sources. This is in stark contrast to a standard surface code with superconducting circuits that would require over a million physical qubits and also in contrast to the all-photonic FBQC architecture using HSPSs studied in this work that we estimate would require roughly 10 billion sources (6 million times more than a RUS-based architecture). Thus a hybrid spin-photonic approach can reduce the resource overhead of fault-tolerant quantum computation by multiple orders-of-magnitude compared to both all-matter and all-photonic platforms. Such a major reduction is expected to translate directly into a substantially reduce cost of fabrication and operation.
Impact of future architecture and component improvements.
This study compares high-performing schemes using best-known approaches, though none of the explored architectures are expected to be optimal and may be improved in the future, perhaps by considering specific hardware details or different resource states.
For instance, it is possible to design photonic schemes using fewer stages of multiplexing, provided that switching networks are highly transmissive but where the majority of loss occurs due to the photonic memory needed to reconfigure the switching network. It may also be possible to determine a more optimal division of the construction into intermediate states or use of auxiliary resources [57].
On a similar note, the 12-source RUS scheme was designed to minimize the optical depth seen by the resource state. However, there are possibly more resource-efficient schemes that combine the caterpillar graph state generation with RUS gates to reach even better upper bounds on resource efficiency. The consequence will be to use at least one stage of photonic multiplexing, which will increase the minimum optical depth. But, in practice, when accounting for the trade-off between resource efficiency and loss tolerance, such a scheme could outperform the 12-source scheme in both metrics in some parameter regimes.
Suitability and limits of current state-of-the-art sources for the proposed RSG architectures.
In addition to the optical depth, it is crucial to consider the efficiency of the source itself. For heralded sources, free-space implementations have shown efficiency values up to 94% [53] before lossy collection into a fiber. On-chip implementations necessary to mass fabricate sources can produce high purity photons, but reaching high source efficiency remains a challenge with values around 26% [54]. Deterministic sources have improved significantly over the past decade, with micropillar cavities consistently demonstrating efficiency values above $50\%$ before lossy collection into a fiber [28, 58]. On-chip sources have reached efficiencies up to 86% [59], but efficient coupling to low-loss waveguides or fibers has yet to be demonstrated. State-of-the-art values are obtained using a tunable open cavity design. This latter approach has demonstrated the highest fibered efficiencies, with $57\%$ in 2020 [52] and $71\%$ in 2023 [30]. However, the manufacturability and stability of open cavity designs remains challenging.
In addition to photon collection efficiency, photons that are highly pure and indistinguishable are necessary to ensure high-fidelity fusion and RUS gates. Currently, HSPS hold the record for indistinguishability of up to 99.5% for photons from distinct sources, but come with a multi-photon error of 0.7% (integrated autocorrelation of $g^{(2)}(0)=0.35\%$) [54]. This also comes along with a low success probability (est. $p_{s}<1\%$) limited by spectral filters that cause $20\%$ loss [54] and hence prevent efficient photon-number resolved heralding that suppresses multi-photon emission produced at higher success probabilities. The state-of-the-art for indistinguishability between photons from distinct QDs in bulk material is 93.0% [60], but is expected to reach values exceeding $99\%$ when using a cavity to suppress phonon side-band emission [61]. Quantum dots also consistently provide sub-1% multi-photon errors without compromising other metrics [58], and they hold the record for multi-photon errors as low as 0.02% [62].
On-chip optical components are also reaching the high levels of quality needed to implement linear-optical processing. Twenty-mode universal SiN chips are commercially available that reach $>97\%$ transformation fidelity [63]. Using machine learning characterization, universal interferometers can also reach fidelity exceeding 99.77% [64]. Smaller specialized integrated circuits show even higher fidelity, with a standard Mach-Zehnder interferometer reaching a fidelity up to 99.996% when characterized with coherent light [65, 54]. Component losses are also improving, with SiN integrated beam splitters and mode crossings reaching 0.012% and 0.037%, respectively [54]. The components that must still be improved are the integrated PNR detectors, the chip-to-fiber interface and high-bandwidth phase shifters. The last two currently incur a loss of 1.3% and 2.3%, respectively [54], which exceed the maximum loss per component estimated for all but the RUS-based architecture.
Solid-state QD sources with a matter qubit degree of freedom have recently demonstrated the generation of small entangled states [36, 37, 38, 39, 17]. These early demonstrations use devices with limited spin coherence times on the order of nanoseconds, which limits graph state sizes. However, it is predicted that 30-photon graph states with $>50\%$ fidelity can be achieved with only minor adjustments [17]. Moreover, state-of-the-art spin coherence times of up to 100 microseconds have been achieved with QDs using dynamical decoupling [66]. This is more than sufficient to construct a 24-photon resource state with $>99.95\%$ fidelity, even when limited by current feedforward times of 23 ns [67]. With this coherence time, resource state fidelity will be primarily limited by single-qubit gates, which can already reach 99.3% [66] and are expected to improve substantially in an industrial setting, in addition to the RUS gate fidelity, which depends primarily on the purity and indistinguishability of the emitted photons.
V Conclusion
The main strengths of photonic quantum computers are modularity, scalability, and flexibility. It is often said that the main drawback is the resource overhead. However, we demonstrated that this overhead practically vanishes when using deterministic single-photon sources with a single matter qubit degree of freedom.
The RUS module architecture builds on these strengths by harnessing a key advantage of material platforms—deterministic entangling gates. This resource-efficient approach allows high-performance sources to be selected and connected to a RUS module, regardless of their location on a chip. Defective sources and components can be easily replaced or upgraded. Additionally, the module can be reconfigured after fabrication, either by rerouting fibers or programming a photonic integrated circuit, to generate different resource states depending on the task.
Our results, combined with recent tailored architectures beyond the fusion-based paradigm leveraging RUS gates and measurements [8, 7, 9], highlight the untapped potential of hybrid platforms. These advancements open critical avenues for developing more efficient fault-tolerant photonic architectures, accelerating progress toward utility-scale quantum computation while minimizing resource costs.
Acknowledgments
The authors thank Grégoire de Gliniasty, Emilio Annoni, and Niccolo Somaschi for fruitful discussions; Rawad Mezher, Nicolas Maring, Jean Senellart, Thomas Volz, and Andrew White for valuable feedback on the manuscript. This work has been co-funded by the European Commission as part of the EIC accelerator program under the grant agreement 190188855 for SEPOQC project and by the Horizon-CL4 program under the grant agreement 101135288 for EPIQUE project.